OpenStudy (goformit100):

A solid cube of aluminum (density 2.7 g/cm3) has a volume of 0.20 cm^3. How many aluminum atoms are contained in the cube?

OpenStudy (anonymous):

find the mass first since density=mass/volume mass = 2.7 x 0.20 = 0.54 g then find the no. of moles no. of moles = mass/molar mass = 0.54/26.98 = 0.020 moles now multiply it by Avogadro's constant no. of molecules = no.of moles x Avogadro's constant = 0.020 x 6.023 x 10^23 = 1.2046 x 10^22 molecules